OpenStudy (callisto):

Coordinate geometry question #4 -Circle

OpenStudy (callisto):

I can reject the options D, and E. But I have some problems with finding the radius. I thought the radius was 5.

OpenStudy (experimentx):

A

OpenStudy (callisto):

I thought so, but the answer is B. That's weird...

OpenStudy (experimentx):

miscalculation .... 4,5 i took height for 4, becuase 4^2+3^2 = 5^5 I can't describe my this type of tendency,

OpenStudy (experimentx):

while the height should have been 5 so r^2 = 5^2+3^2

OpenStudy (callisto):

Why is the height 5?

OpenStudy (callisto):

Oh..Nothing, I read the question wrong. Thanks

OpenStudy (experimentx):



OpenStudy (callisto):

Thanks :)

OpenStudy (experimentx):

yw